Appendix B Installation Procedures – PC and Generic

B.1. Introduction

The AMS-TEX software can be used with any full implementation of TEX. Sections B.2–B.5 of this appendix describe the installation of AMS-TEX for DOS implementations of TEX on an IBM PC or compatible machine from a DOS diskette or diskettes provided by the American Mathematical Society. PCTEX (from Personal TEX, Inc.) and emTEX are used as examples. If your operating system is not DOS, or if you obtained AMS-TEX through other channels, you won’t be able to use the install program that is provided. Instead, follow the generic installation instructions in section B.6. You will have to refer to the documentation for your particular implementation of TEX to work out the details of any necessary variations. Users who encounter difficulties should seek help from the manufacturer of their implementation of TEX, or from the American Mathematical Society’s technical support staff.

The TFM files for some of AMSFonts 2.2 (msam*, msbm*, and eufm*) are needed to run AMS-TEX with the AMSPPT documentstyle, even if you don’t plan to actually print anything using the AMSFonts. (These fonts are needed to process and print the User’s Guide of which this index is a part.) In the diskette distribution TFM files for all AMSFonts are provided in a directory named \tfm. They are also available from the AMS Internet archive, e-math.ams.org, by anonymous FTP.

B.2. Getting Started

The installation procedure consists of moving files to the proper directories, and creating format files. The first step is to determine the names of the proper directories, which will vary with different implementations of TEX. You need to know the name of your TEX inputs directory, your TEX TFM directory, and your TEX formats directory. Check your directory structure or consult the documentation for your implementation of TEX to see what these names are. If you have any difficulty determining these directories:

(a) Search for the plain TEX input file, plain.tex. The directory where it is located will be your TEX inputs directory.

(b) Search for the font file cmr10.tfm. The directory where it is found will be your TEX TFM directory.

(c) Search for the plain TEX format file, plain.fmt. The directory where it is found will be your TEX formats directory.

For example, for PCTEX the directory names are \pctex\texinput,

\pctex\textfms, and \pctex\texfmts.

Note: If you currently have any of the following files from earlier releases of AMS-TEX, backup and delete them before installing the new version. They are either irrelevant or superseded in the new version of AMS-TEX, and it is best to remove them to avoid confusion. All of them except for the last one would be found in your TEX inputs directory; amsplain.fmt would be found in your TEX formats directory.

  amsfil.chg  amsplain.tex  amsppt.sty    amsppt.mor    amstex.tex
  amstex.chg  cyracc.def    cyrmemo.def   cyrmemo.tex   amsplain.fmt

Users who did not receive AMS-TEX on diskette from the AMS, or those using non-DOS implementations of TEX, should proceed now to section B.6.

B.3. Installing AMS-TEX (DOS/diskette)

The following files are used in the installation process for DOS implementations of TEX, for AMS-TEX 2.1:

    install.exe            amsinst.bat            amstex.ini

For DOS installation, you need to know on which drive to install AMS-TEX. This should be the drive on which you already have installed TEX.

For PCTEX, you will probably want to select the first choice for each of the questions which the installation programs asks you. This will result in placing the AMS-TEX files in the directory \pctex\texinput and the TFM files in the

directory \pctex\textfms.

For other DOS implementations of TEX, you will need to know the names of your TEX inputs directory and TEX TFM directory, as described earlier, so that you can enter them when prompted.

When you have the disk and directory information ready, place the disk labeled “AMS-TEX 2.1” in your floppy disk drive and type the following commands (if the disk is in drive B, substitute “b” for “a” in the first line):

a: ⟨return⟩ install ⟨return⟩

This will run the install program, which will ask you some questions. Enter the answers which you determined from reading the above paragraphs.

B.4. Running INITEX to Create Format Files (DOS)

To complete the installation procedure, you should create a format file. This will enable you to run TEX with AMS-TEX, or AMS-TEX and the AMSPPT preprint style, preloaded. This preloading will save quite a bit of startup time on slower systems. Note: Each format file takes up 150K–300K of disk space (depending on your implementation of TEX).

Warning: INITEX requires much more memory to run than regular TEX. The first time you try to run INITEX, you may get the message “Not enough memory to run TeX” or something similar (with simpler implementations, it may just crash). If this happens, you must remove as many memory-resident programs as possible (such as communications software and memory-resident utilities [TSRs]) and reboot your system to create enough memory to run INITEX. Consult the documentation for your implementation of TEX for more information on running INITEX.

You are now ready to create a format file so that AMS-TEX, or AMS-TEX and the AMSPPT preprint style, can be preloaded when you typeset a document. Before creating your format file, you will want to consider whether you habitually use the AMSPPT documentstyle. If you use other documentstyles rarely or never, then you would benefit from the use of a format file with amsppt.sty preloaded. If you are likely to use other documentstyles periodically, then you probably do not want to preload amsppt.sty. To make a simple AMS-TEX format file, proceed with the next paragraph. To make a format file with amsppt.sty preloaded, edit the file amstex.ini and remove the percent sign (comment character) at the beginning of the line \documentstyle{amsppt}, just before the \dump command. If you intend to use Type-1 outline versions of AMSFonts, read Appendix D before proceeding.

The TEX file named amstex.ini should now be installed in your TEX inputs directory. Once you have decided whether to create an AMS-TEX or an AMSPPT format file (see previous paragraph), run INITEX on amstex.ini: For PCTEX, give the command

     tex amstex.ini -i  ⟨return⟩

For emtex, give the command

     tex386 -i amstex.ini  ⟨return⟩

For other implementations of TEX the form of the INITEX command may be

different, e.g., tex/i amstex.ini.

This creates an AMS-TEX format file named amstex.fmt. If you included AMSPPT, you may wish to rename it amsppt.fmt.

Move the format files to the right directory. Some implementations of TEX, including PCTEX, will automatically place the format file in the proper directory. Otherwise you should now move the file manually into your TEX formats directory.

B.5. Using AMS-TEX 2.0+ (DOS)

On the distribution diskettes, or in a distribution obtained by FTP or from a Web site, a DOS batch file amstex.bat is provided, to make use of the format file more convenient. If you received the distribution on diskette, amstex.bat will have been placed by the installation procedures in the root directory of the drive where TEX is located; otherwise, you should install it there manually. Then, if you are connected to that directory or if it is in your system path, you would run AMS-TEX on a file called filename.tex by typing

     amstex filename  ⟨return⟩

B.6. Installing AMS-TEX (Generic)

Use these instructions if you didn’t receive AMS-TEX on diskettes from the AMS, but obtained it by other methods, or if you have a non-DOS implementation of TEX. You will need to know the names of your TEX inputs directory, TEX TFM directory, and TEX formats directory, as explained in section B.2.

(1) Copy the following files into your TEX inputs directory: amstex.tex,

       amsppt.sty, amssym.tex, amstex.ini, and amsppt1.tex.

(2) There are three documentation files: amsguide.tex, joyerr.tex, and amsppt.doc. If you have a documentation directory, or wish to create one, then put these files there; otherwise they can be put in the TEX inputs directory.

(3) Copy the AMSFonts TFM files to your TEX TFM directory. If you don’t plan to use any of the AMSFonts, you should still copy dummy.tfm, which is needed for AMS-TEX’s syntax check option, and msam*.tfm, msbm*.tfm, and eufm*.tfm, which are needed for typesetting the User’s Guide.

(4) You are now ready to create a format file so that AMS-TEX, or AMS-TEX and the AMSPPT preprint style, can be preloaded when you typeset a document. Before creating your format file, you will want to consider whether you habitually use the AMSPPT documentstyle. If you use other documentstyles rarely or never, then you would benefit from the use of a format file with amsppt.sty preloaded. If you are likely to use other documentstyles periodically, then you probably do not want to preload amsppt.sty. To make a simple AMS-TEX format file, proceed with the next paragraph. To make a format file with amsppt.sty preloaded, edit the file amstex.ini and remove the percent sign (comment character) at the beginning of the line \documentstyle{amsppt}, just before the \dump command. If you intend to use Type-1 outline versions of AMSFonts, read Appendix D before proceeding.

(5) Check the documentation for your implementation of TEX to find out how to run INITEX and create format files. Format files greatly speed up processing when you are using a large macro package such as AMS-TEX. If your implementation of TEX doesn’t automatically place format files in the TEX formats directory (check your documentation), you will have to either go to the TEX formats directory before running INITEX, or move the format files there after they are created.

If you are a PC user, return to section B.4, read the warning, then proceed to

the paragraph beginning “The TEX file named amstex.ini”. Run the file amstex.ini through INITEX, to create the file amstex.fmt, which

is a preloaded form of AMS-TEX. If you do not have an executable program or symbol initex, a typical command to perform this function is

  tex -i amstex.ini  ⟨return⟩

You will now have a file amstex.fmt. If you included AMSPPT, you may wish

to rename it amsppt.fmt.

(6) Move the format file that you just created into your TEX formats directory, if you didn’t create it there.

For details of how to use format files with your implementation of TEX, see your documentation. Typically, you use a format file by specifying it on the command line preceded by an ampersand, e.g., tex &amstex filename.

B.7. Getting a printed copy of the User’s Guide

If you did not receive a printed copy of this User’s Guide with your distribution, you can use your newly created AMS-TEX or AMSPPT format file to typeset the

file amsguide.tex.

Appendix C Installation Procedures – Macintosh (Textures)

C.1. Introduction

These instructions describe the installation of AMS-TEX for use with Textures, on the Macintosh. There is one disk in the distribution. The files amstex.tex and amsppt.sty are Textures documents to be input by Textures. The file amsppt.doc is technical documentation for the macros defined in amsppt.sty. The file amstex.ini is used in the installation to create format files. The files amsguide.tex and amstinst.tex are the Textures source for this User’s Guide.

The file joyerr.tex is a list of errata to the 1986 edition of The Joy of TEX. You may typeset it in Textures using your new AMS-TEX format file (once it has been created). If you do not have AMSFonts 2.2, the characters in this file which come from the AMS symbol fonts will appear in the default system font. The file AMSFonts 2.2 metrics contains the TEX metrics information for AMSFonts 2.2. This file is required to use amsppt.sty even if you do not have AMSFonts 2.2. (Note: Even if you currently have AMSFonts 2.0, we strongly recommend that you obtain version 2.2, as the metrics have changed.)

AMS-TEX Versions 2.0+ can be used with Textures without AMSFonts. However, AMS-TEX Versions 2.0+ will not work with releases of AMSFonts previous to Version 2.0. Additionally, if AMSFonts 2.2 are to be used, they require Textures version 1.2 or higher. If you have an earlier release of Textures, you must upgrade before using these fonts. To upgrade Textures, contact the manufacturer: Blue Sky Research (see below, Getting Help).

Note: AMSFonts are available in Type-1 outline form from Blue Sky Research. If you are using these fonts, read Appendix D before proceeding.

C.2. If You Have a Previous Version of AMS-TEX

Before installing the new version of AMS-TEX, you should backup and delete from your hard disk any files which are related to previous releases of AMS-TEX. In particular, you should remove the following files from your Textures folder (or any folders contained in your Textures folder):

  amsfil.chg  amsplain.tex  amsppt.sty    amsppt.mor    amstex.tex
  amstex.chg  cyracc.def    cyrmemo.def   cyrmemo.tex

and you should remove any previous AMS-TEX format files from your TeX formats folder. These files are either irrelevant or superseded in the new version of AMS-TEX, and it is best to remove them from your hard disk to avoid confusion.

C.3. Installing AMS-TEX

If you do not have a folder named TeX inputs inside the Textures folder on your hard disk, create a new folder inside your Textures folder and name it TeX inputs. Also, if you do not have a TeX formats folder, create a new folder inside your Textures folder and name it TeX formats. Likewise, create a TeX fonts folder if you do not already have one.

To install the new version of AMS-TEX on your system, copy the following files into the TeX inputs folder inside the Textures folder on your hard disk:

      amstex.tex  amssym.tex  amsppt.sty  amsppt1.tex  amstex.ini

Then copy the file amsfonts 2.2 metrics into the TeX fonts folder inside of the Textures folder on your hard disk.

You are now ready to create a format file so that AMS-TEX, or AMS-TEX and the AMSPPT preprint style, can be preloaded when you typeset a document. Before creating your format file, you will want to consider whether you habitually use the AMSPPT documentstyle. If you use other documentstyles rarely or never, then you would benefit from the use of a format file with amsppt.sty preloaded. If you are likely to use other documentstyles periodically, then you probably do not want to preload amsppt.sty. To make a simple AMS-TEX format file, proceed with the next paragraph. To make a format file with amsppt.sty preloaded, edit the file amstex.ini and remove the percent sign (comment character) at the beginning of the line \documentstyle{amsppt}, just before the \dump command. If you intend to use Type-1 outline versions of AMSFonts, read Appendix D before proceeding.

Make sure that Flash mode is off. You cannot create a format with Flash mode enabled.

Open the TeX inputs folder (inside your Textures folder) and double-click on the file amstex.ini to start running Textures. Make sure that the line \input plain is commented out (begins with a percent sign). If there is not a percent sign at the beginning of this line, insert one. Check under the Typeset menu to make sure that the Plain format is selected (there is a check mark next to it). Select “Typeset” from the Typeset menu. When the dialog box appears asking you what to name the format file, go through the folder hierarchy to place yourself inside the TeX formats folder inside your Textures folder. You may name the file what you wish, but 2.1 is a good idea. If you included amsppt.sty in your format file, name the format file AMS-TeX/AMSPPT or something similarly descriptive.

For more information about using formats in Textures, see the Textures User’s Guide.

If you did not receive a printed copy of the AMS-TEX User’s Guide with your distribution, you can use your newly created AMS-TEX or AMSPPT format file to typeset the file amsguide.tex.

C.4. Getting Help

Questions concerning Textures should be directed to:

Blue Sky Research

317 Southwest Alder, #1200

Portland, OR 97204

Phone: 800-622-8398 or 503-222-9571

Appendix D Before Installing AMS-TEX: Facts About Fonts

D.1. Using AMSFonts in PostScript Type-1 form

The AMSFonts Version 2.2 have been converted to PostScript Type-1 outlines in two forms:

Blue Sky Research and Y&Y have created a collection containing all the typefaces, but only in 5, 7, and 10 point sizes; other sizes must be created by scaling. The AMS-TEX preprint style provides a way for authors to indicate that fonts should be loaded under this alternative scaling convention. If you are using the Blue Sky/Y&Y version of AMSFonts, insert the line \PSAMSFonts in the preamble of your file, before the \topmatter line.

Basil Malyshev has created a collection, called BaKoMa, containing the Computer Modern fonts as well as all sizes of the AMSFonts used in mathematics, but excluding the wncy* cyrillic fonts. The BaKoMa fonts can be used with no special action.

D.2. Using AMS Symbol Fonts

The preprint style automatically loads the Fraktur font and both fonts of extra symbols (msam and msbm), as well as all the symbol names, as described in the sections Fonts and Symbol Names. If these will not be used, and you want a version of amsppt.sty that requires less memory than the default version, you can suppress the loading of these resources. If amsppt.sty is to be included in your format file, you must make this change before creating the format file.

An “initialization file”, amsppt.sti, contains settings for some options that a user may wish to change locally. The following two lines appear in this file:

        \loadeufm \loadmsam \loadmsbm
        \message{symbol names}\UseAMSsymbols\message{,}

To use amsppt.sty without AMSFonts, comment out both lines (place a % at the beginning of each line); to disable just the symbol names, comment out the second line. Extra math symbols can be loaded on demand with \newsymbol or by including \UseAMSsymbols in the preamble of a particular document.